For a circle of radius 9 feet, find the arc length s cut off by a central angle of 6°.
1
Expert's answer
2012-08-21T09:29:28-0400
f the measure of central angle is 6°, the arc measure is 2*6°=12°.
Then use this proportion:
12° - x
360° - 2*pi*R
x = 2*pi*R * 12° / 360° = 2*9*3.14 / 30 = 1.88 feet.